# Can You Hear Me Now? A Benchmark for Long-Range Graph Propagation

### v1

@@ 5 Conclusion
In this paper we propose ECHO, a new benchmark for evaluating long-range information propagation in GNNs. Our benchmark includes two main components, ECHO-Synth and a set of chemically grounded real-world datasets (collectively referred to as ECHO-Chem), that target long-range communication in both synthetic and real-world settings. The synthetic tasks are designed to predict algorithmic and long-range-by-design graph properties, while the real-world tasks focus on predicting atomic charge distributions and molecular total energies, both of which critically depend on long-range quantum interactions. We provided a detailed analysis to demonstrate that the tasks in ECHO genuinely capture long-range dependencies, and we established strong baselines for each task to provide a comprehensive reference point for future research. Our results highlight the limitations of current GNN architectures when faced with long-range propagation challenges, and we believe that ECHO will serve as a critical step toward building more robust, scalable, and generalizable GNNs capable of handling the full spectrum of graph-based learning tasks, posing a challenge to the community to push the boundaries of GNN design and evaluation. Not only does ECHO provide a solid benchmark, but it also leaves ample room for future architectures to improve and advance GNN architectures capable of more effective information propagation.
Reproducibility
@@
We openly release data and code to replicate our results and perform new evaluations at https://github.com/Graph-ECHO-Benchmark/ECHO.
Appendix A Discussion on LRGB
@@
One of the most widely used benchmark for assessing the lon-range propagation capabilities of GNNs is the Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022). The benchmark proposes five tasks: two molecular property prediction tasks (Peptides-func and Peptides-struct), one molecular bond prediction task (PCQM-Contact), and two computer vision tasks (PascalVOC-SP and COCO-SP). However, despite initial rapid improvements, performance on LRGB has plateaued. Since its introduction in 2022, there has been a noticeable deceleration in progress. Considering a set of 30 models on the Peptides-func task, we observe a performance improvement by 6.5% in the first year, but only by 1.3% in the second, and no significant gain in the third year (Gravina et al., 2025; Heilig et al., 2025; Errica et al., 2024; Gutteridge et al., 2023; Dwivedi et al., 2022; Tönshoff et al., 2023; Giusti et al., 2023; Ma et al., 2023a; Shirzad et al., 2023; Behrouz & Hashemi, 2024; Wang et al., 2024; Eliasof et al., 2025; Ding et al., 2024; Ma et al., 2023b; Cai et al., 2023; Glickman & Yahav, 2023; He et al., 2023; Rampášek et al., 2022; Ngo et al., 2023; Michel et al., 2023; Geisler et al., 2024; Choi et al., 2024). A similar trend exists for the other benchmark tasks as well.
@@
Furthermore, a recent analysis on LRGB (Bamberger et al., 2025b), as well as the benchmark’s sensitivity to hyperparameter tuning (Tönshoff et al., 2023), raises additional concerns about the long-range nature of its tasks. The analysis reveals that only a subset of tasks genuinely require longer interactions, while the peptides tasks are effectively local. This highlights the need for more focused benchmarks that explicitly and systematically test long-range propagation capabilities of GNNs.
Appendix B Chemical Simulation Technical Information
@@
3D Structure Generation from SMILES. Since subsequent charge optimization steps require pre-optimized 3D coordinates, all structures were geometry-optimized prior to simulation using Open Babel (O’Boyle et al., 2011) and its Python interface, Pybel (O’Boyle et al., 2008) Initial molecular geometries were generated from SMILES strings using the General AMBER Force Field (GAFF) (Grimme et al., 2010). GAFF was chosen over alternatives such as MMFF94 (Halgren, 1996) due to its favorable trade-off between accuracy and computational cost, and its strong performance in predicting both energies and geometries. The optimization procedure involved 100 steps of coarse minimization followed by 500 steps of local refinement for each molecule. The SMILES strings were converted into 3D conformers, which were then minimized to yield low-energy structures. These structures were exported in SDF format for subsequent compatibility. The average time required for 3D structure generation per molecule—considering only those satisfying the ECHO-Charge and ECHO-Energy dataset diameter criteria—was $\mathbf{562\pm 124}$ ms.
Quantum Chemical Computations with ORCA. To compute partial atomic charges and total energies, we employed the ORCA quantum chemistry software suite (version 6.0.1) (Neese, 2022; 2023; Neese et al., 2020). All calculations were performed using the B3LYP a hybrid density functional (DFT) method that mixes Hartree–Fock exchange with Becke’s exchange and Lee–Yang–Parr correlation functionals to balance accuracy and efficiency in quantum chemical calculations (Argaman & Makov, 2000).
Table 3: Mean time required for computation of partial charges of a single molecule with different configuration of the ORCA tool. Variance is computed over 30 random molecules from the ECHO-Charge/ECHO-Energy dataset. Denotes the chosen basis set for DFT computation.
@@
We provide a summary of required times for computation with both full DFT computations and HF methods in Table 3. Under our configuration, the average runtime for a single quantum chemical calculation was $\mathbf{634.5\pm 21.3}$ seconds per molecule requiring $\approx$ 2 months of computational time on our hardware configuration. Simulation were run exploiting full thread parallelism provided by ORCA.
@@
Charge Extraction. Atomic partial charges were extracted using the Löwdin Szabo & Ostlund (1989) population analysis method. These charges were used as supervision signals in our dataset generation pipeline.
Appendix C Hardware Resources
@@
Appendix D Additional Dataset Information
@@
Figure 7: Correlation between number of nodes and graph diameter in ECHO-Synth, ECHO-Charge and ECHO-Energy.
Appendix E Hyperparameter Selection
@@
Appendix F Additional Experimental Analysis
In this section, we investigate how the radius of the explored neighborhood influences the performance of each method, as well as how the models perform across graphs with varying diameters. We also analyze how the performance of the models change across different graph topologies and the influence of the readout depth in ECHO-Synth.
F.1 Layer-wise Performance Analysis
@@
  • On the diam task (Figure 8(a)), performance trends are model-dependent. Long-range models such as DRew and A-DGN remain stable or slightly improve, others like PH-DGN exhibit a large performance improvement moving from shallow to medium depth regime. This task, being graph-level and heavily reliant on global information by design, clearly benefits from increased depth and non-dissipative architectures which are able to perform many message-passing steps across multiple hops.
@@
  • Finally, on the ECHO-Charge task (Figure 8(d)), the behavior differs. This task involves precise regression of atomic partial charges, where small errors matter. Most models show stable MAE across depths, except for GCN and GIN, which degrade significantly in the deep regime. Importantly, models with explicit long-range message-passing capabilities (A-DGN, SWAN, PH-DGN, GPS, and DRew) retain high accuracy even at >17>17 layers. This suggests their robustness in fine-grained, long-range molecular prediction tasks. We do not include ECHO-Energy in this ablation, as it exhibited very similar behavior to ECHO-Charge.
Overall, the observed patterns reveal a clear correlation between the number of message-passing layers and performance: models require many layers to perform well, confirming the long-range nature of these benchmarks. Remarkably, architectures explicitly designed to support many message-passing steps consistently outperform others, further confirming the long-range nature of our proposed benchmarks.
@@
F.2 Performance Across Graph Diameters
@@
For the diam task (Figure 9(a)), most models show robust performance for small to moderate diameters, with a slight increase in MAE for very large diameters. Notably, GCN, GraphCON and GCNII architectures exhibits substantial degradation as diameter increases, suggesting poor scalability in capturing global structure on many message-passing steps. Again, non-dissipative architectures (i.e., A-DGN, PH-DGN, and SWAN), DRew and GPS remain consistently accurate across all graph diameters, demonstrating their capacity to generalize across different graph scales. Finally, we observe that models with lower generalization capabilities tend to bias their predictions towards the statistical mean of the dataset. Consequently, these models achieve low error rates in the central range, where the ground truth aligns with the mean, but show high errors at the tails, as they fail to distinguish graphs with extreme diameters from the average case.
@@
The ecc task (Figure 9(b)) reveals a characteristic U-shaped curve. Performance improves as diameter increases from small to moderate values, and deteriorates again for very large graphs. Here, all models follow a similar trend, although A-DGN and GPS tend to dominate in the optimal range. This patterns is highly correlated with the node eccentricity distribution depicted in Figure 4, as it follows a non-uniform distribution, creating an inverse relationship between sample frequency and error. In particular, in the center of the plot, it is possible to see the GNNs that learn the dominant pattern; in addition, simply predicting values close to the statistical mean minimizes the expected loss in this dense region, and results in a lower error. Conversely, graph exhibiting lower eccentricities are statistically less frequent, leading to a noticeable degradation of the performance across extreme values.
@@
In the sssp task (Figure 9(c)), increasing graph diameter consistently correlates with rising MAE. Model performance divides into three groups, with GraphCON exhibiting the worst performance both in terms of overall MAE and degradation with increasing diameter. GCN, GCNII, and GIN show similar values across the task and similar degradation trends. Finally, non-dissipative models, GPS Transformer and DRew, once again demonstrate remarkable and consistent performance even on difficult graphs. This trend reinforces the long-range nature of the task, where deeper or more expressive models are required to maintain strong performance.
On the molecular ECHO-Charge task (Figure 9(d)), test MAE consistent across all ranges, but subtle trends emerge. Models like DRew and GPS show stability and even slight improvements for larger molecular graphs, while GCN and GIN degrade more noticeably, confirming their limited capacity to manage increasing molecular complexity. Interestingly, GINE performs substantially better than its counterpart GIN, suggesting that edge-level attributes play a crucial role in the ECHO-Charge regression task. Similarly to the previous ablation study, we do not include ECHO-Energy, as it exhibited very similar behavior to ECHO-Charge. Additionally, we note that all models exhibit a general performance drop when processing molecular graphs with a diameter greater than 35. We attribute this behavior to the original ChEMBL dataset’s distribution, which includes fewer graphs with diameters in the 35–40 range. This also impacts our ECHO-Charge dataset as illustrated in Figure 5. As a result, models have limited opportunity to learn effective representations for such large graphs, which likely contributes to the observed degradation.
@@
F.3 Impact of readout layer depth
@@
F.4 Performance across different graph topologies
In this section we analyze the models’ performance on different graph topologies in ECHO-Synth. Figures 10, 11 and 12 show respectively diam, ecc and sssp. The results show that although absolute performance varies slightly with the underlying graph topology (e.g., GPS performs better on lobster graphs than on tree-like topologies in SSSP), the relative ranking of models remains consistent, reinforcing the robustness of our findings. Interestingly, we observe that the line topology is consistently the most challenging one. This is expected: in a line graph, every message must pass through a sequence of intermediate nodes, making each node a critical bottleneck for information propagation and amplifying the need for long-range communication.
@@ F.4 Performance across different graph topologies
Appendix G Extended Results
In Table 18 we report additional result on ECHO-Synth benchmark. In particular we report MAE, MSE and loss (defined as $\log_{10}(\mathrm{MSE})$) obtained on the test set. Similarly, we report the same metrics for ECHO-Charge and ECHO-Energy in Table 19 and Table 20, respectively.
@@ Appendix H Runtimes
Metric Model diam sssp ecc ECHO-Charge ECHO-Energy
@@
Training (s) PH-DGN $2.840_{\pm 0.060}$ $4.480_{\pm 0.060}$ $3.010_{\pm 0.060}$ $39.405_{\pm 1.753}$ $49.099_{\pm 1.186}$
@@
Appendix I Visualization of GPS Attention Patterns

### v2

@@ 5 Conclusion
In this paper we propose ECHO, a new benchmark for evaluating long-range information propagation in GNNs. Our benchmark includes two main components, ECHO-Synth and a set of chemically grounded real-world datasets (collectively referred to as ECHO-Chem), that target long-range communication in both synthetic and real-world settings. The synthetic tasks are designed to predict algorithmic and long-range-by-design graph properties, while the real-world tasks focus on predicting atomic charge distributions and molecular total energies, both of which critically depend on long-range quantum interactions. We provided a detailed analysis to demonstrate that the tasks in ECHO genuinely capture long-range dependencies, and we established strong baselines for each task to provide a comprehensive reference point for future research. Our results highlight the limitations of current GNN architectures when faced with long-range propagation challenges, and we believe that ECHO will serve as a critical step toward building more robust, scalable, and generalizable GNNs capable of handling the full spectrum of graph-based learning tasks, posing a challenge to the community to push the boundaries of GNN design and evaluation. Not only ECHO provides a solid benchmark, but it also leaves ample room for future architectures to improve and advance GNN architectures capable of more effective information propagation.
Ethics Statement
The research conducted in this paper conforms in every aspect with the ICLR Code of Ethics. Our study does not involve human subjects, sensitive personal data, or applications with foreseeable harmful consequences. No ethical concerns are anticipated regarding data usage, methodology, or findings.
Reproducibility Statement
@@
We provide all necessary details to reproduce our ECHO benchmark in Section 3 and Appendix C, and describe the setup of each experiment in Section 4 and Appendix F, thus ensuring sufficient information to replicate our results. We openly release data and code to replicate our results and perform new evaluations at https://github.com/Graph-ECHO-Benchmark/ECHO.
Acknowledgments
This work has been partially supported by EU-EIC EMERGE (Grant No. 101070918) and NextGeneration EU programme PNNR Partenariato Esteso PE00000013 “FAIR” - Future Artificial Intelligence Research” - Spoke 1 “Human-centered AI”.
Appendix A LLMs Usage
Large Language Models (LLMs) were used as general-purpose assistive tools to improve the writing quality of this paper. Specifically, we used LLMs to help with grammar correction, rephrasing for clarity, and suggesting some improvements to the overall structure of the text. All LLM-generated text was carefully reviewed and edited by the authors to ensure that it accurately reflects the authors’ intentions and scientific content. No LLMs were used to generate scientific content, including but not limited to research direction, hypothesis formulation, experimental design, data analysis, or interpretation of results.
Appendix B Discussion on LRGB
@@
One of the most widely used benchmarks for assessing the long-range propagation capabilities of GNNs is the Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022). The benchmark proposes five tasks: two molecular property prediction tasks (Peptides-func and Peptides-struct), one molecular bond prediction task (PCQM-Contact), and two computer vision tasks (PascalVOC-SP and COCO-SP). However, despite initial rapid improvements, performance on LRGB has plateaued. Since its introduction in 2022, there has been a noticeable deceleration in progress. Considering a set of 30 models on the Peptides-func task, we observe a performance improvement of 6.5% in the first year, but only by 1.3% in the second, and no significant gain in the third year (Dwivedi et al., 2022; Rampášek et al., 2022; Tönshoff et al., 2023; Shirzad et al., 2023; Cai et al., 2023; Glickman & Yahav, 2023; He et al., 2023; Gutteridge et al., 2023; Giusti et al., 2023; Ma et al., 2023a; Ngo et al., 2023; Michel et al., 2023; Ma et al., 2023b; Ding et al., 2024; Choi et al., 2024; Geisler et al., 2024; Behrouz & Hashemi, 2024; Wang et al., 2024; Errica et al., 2024; Gravina et al., 2025; Heilig et al., 2025; Eliasof et al., 2025; Hariri et al., 2025; Trenta et al., 2025). A similar trend exists for the other benchmark tasks as well.
@@
Furthermore, a recent analysis on LRGB (Bamberger et al., 2025b), as well as the benchmark’s sensitivity to hyperparameter tuning (Tönshoff et al., 2023), raises additional concerns about the long-range nature of its tasks. The analysis reveals that only a subset of tasks genuinely require longer interactions, while the peptides tasks are effectively local. This highlights the need for more focused benchmarks that explicitly and systematically test the long-range propagation capabilities of GNNs.
Appendix C Chemical Simulation Technical Information
@@
3D Structure Generation from SMILES. Since subsequent charge optimization steps require pre-optimized 3D coordinates, all structures were geometry-optimized prior to simulation using Open Babel (O’Boyle et al., 2011) and its Python interface, Pybel (O’Boyle et al., 2008) Initial molecular geometries were generated from SMILES strings using the General AMBER Force Field (GAFF) (Grimme et al., 2010). GAFF was chosen over alternatives such as MMFF94 (Halgren, 1996) due to its favorable trade-off between accuracy and computational cost, and its strong performance in predicting both energies and geometries. The optimization procedure involved 100 steps of coarse minimization followed by 500 steps of local refinement for each molecule. The SMILES strings were converted into 3D conformers, which were then minimized to yield low-energy structures. These structures were exported in SDF format for subsequent compatibility. The average time required for 3D structure generation per molecule (considering only those satisfying the ECHO-Charge and ECHO-Energy dataset diameter criteria) was $\mathbf{562\pm 124}$ ms.
Quantum Chemical Computations with ORCA. To compute partial atomic charges and total energies, we employed the ORCA quantum chemistry software suite (version 6.0.1) (Neese, 2022; 2023; Neese et al., 2020). All calculations were performed using the B3LYP, a hybrid density functional (DFT) method that mixes Hartree–Fock exchange with Becke’s exchange and Lee–Yang–Parr correlation functionals to balance accuracy and efficiency in quantum chemical calculations (Argaman & Makov, 2000).
Table 3: Mean time required for computation of partial charges of a single molecule with different configurations of the ORCA tool. Variance is computed over 30 random molecules from the ECHO-Charge/ECHO-Energy dataset. Denotes the chosen basis set for DFT computation.
@@
We provide a summary of required times for computation with both full DFT computations and HF methods in Table 3. Under our configuration, the average runtime for a single quantum chemical calculation was $\mathbf{634.5\pm 21.3}$ seconds per molecule, requiring $\approx$ 2 months of computational time on our hardware configuration. Simulations were run exploiting full thread parallelism provided by ORCA.
@@
Charge Extraction. Atomic partial charges were extracted using the Löwdin population analysis method (Szabo & Ostlund, 1989). These charges were used as supervision signals in our dataset generation pipeline.
Appendix D Hardware Resources
@@
Appendix E Additional Dataset Information
@@
Figure 7: Correlation between the number of nodes and graph diameter in ECHO-Synth, ECHO-Charge and ECHO-Energy.
Appendix F Hyperparameter Selection
@@
Appendix G Additional Experimental Analysis
In this section, we investigate how the radius of the explored neighborhood influences the performance of each method, as well as how the models perform across graphs with varying diameters. We also analyze how the performance of the models changes across different graph topologies and the influence of the readout depth in ECHO-Synth.
G.1 Layer-wise Performance Analysis
@@
  • On the diam task (Figure 8(a)), performance trends are model-dependent. Long-range models such as DRew and A-DGN remain stable or slightly improve, and PH-DGN exhibits a large performance improvement moving from a shallow to a medium depth regime. This task, being graph-level and heavily reliant on global information by design, clearly benefits from increased depth and non-dissipative architectures which are able to perform many message-passing steps across multiple hops.
@@
  • On the ECHO-Charge task (Figure 8(d)), the behavior differs. This task involves precise regression of atomic partial charges, where small errors matter. Most models show stable MAE across depths, except for GCN and GIN, which degrade significantly in the deep regime. Importantly, models with explicit long-range message-passing capabilities (A-DGN, SWAN, PH-DGN, and GPS) retain high accuracy even at >17>17 layers. This suggests their robustness in fine-grained, long-range molecular prediction tasks.
  • Finally, on the ECHO-Energy task (Figure 8(e)), performance trends are more model-dependent and exhibit higher variability, while still underlying the benefit of deeper regimes. Generally, non-dissipative architectures such as PH-DGN and A-DGN benefit from increased depth. The best performance is achieved by the transformer-based model GPS, supporting the intuition that this graph-level task requires globally informed representations.
Overall, the observed patterns reveal a clear correlation between the number of message-passing layers and performance: models require many layers to perform well, confirming that these benchmarks require information to propagate over long distances. Remarkably, architectures explicitly designed to support many message-passing steps consistently outperform others, further confirming the long-range nature of our proposed benchmarks.
@@
G.2 Performance Across Graph Diameters
@@
For the diam task (Figure 9(a)), most models show robust performance for small to moderate diameters, with a slight increase in MAE for very large diameters. Notably, GCN, GraphCON and GCNII architectures exhibit substantial degradation as diameter increases, suggesting poor scalability in capturing global structure on many message-passing steps. Again, non-dissipative architectures (i.e., A-DGN, PH-DGN, and SWAN), DRew, GPS, and GRIT remain consistently accurate across all graph diameters, demonstrating their capacity to generalize across different graph scales. Finally, we observe that models with lower generalization capabilities tend to bias their predictions towards the statistical mean of the dataset. Consequently, these models achieve low error rates in the central range, where the ground truth aligns with the mean, but show high errors at the tails, as they fail to distinguish graphs with extreme diameters from the average case.
@@
The ecc task (Figure 9(b)) reveals a characteristic U-shaped curve. Performance improves as diameter increases from small to moderate values, and deteriorates again for very large graphs. Here, all models follow a similar performance pattern that is highly correlated with the node eccentricity distribution depicted in Figure 4, creating an inverse relationship between sample frequency and error. In particular, in the center of the plot, it is possible to see the GNNs that learn the dominant pattern; in addition, simply predicting values close to the statistical mean minimizes the expected loss in this dense region, and results in a lower error. Conversely, graphs exhibiting lower eccentricities are statistically less frequent, leading to a noticeable degradation of the performance across extreme values.
@@
In the sssp task (Figure 9(c)), increasing graph diameter consistently correlates with rising MAE. Model performance divides into three groups, with GraphCON exhibiting the worst performance both in terms of overall MAE and degradation with increasing diameter. GCN, GCNII, and GIN show similar values across the task and similar degradation trends. Finally, non-dissipative models, Graph Transformers, and DRew once again demonstrate remarkable and consistent performance even on difficult graphs. This trend reinforces the long-range nature of the task, where deeper or more expressive models are required to maintain strong performance.
On the molecular ECHO-Charge task (Figure 9(d)), test MAE consistent across all ranges, but subtle trends emerge. Models like DRew and GPS show stability and even slight improvements for larger molecular graphs. Interestingly, GINE performs substantially better than its counterpart GIN, suggesting that edge-level attributes play a crucial role in the ECHO-Charge regression task.
On the molecular task ECHO-Energy (Figure 9(e)), a clear trend emerges: test MAE increases with graph diameter across all models. However, architectures such as SWAN and DRew exhibit substantially lower sensitivity to diameter, maintaining a nearly constant MAE across the full range. Additionally, we note that all models exhibit a general performance drop when processing molecular graphs with a diameter greater than 35. We attribute this behavior to the original ChEMBL dataset’s distribution, which includes fewer graphs with diameters in the 35–40 range, as illustrated in Figure 6.
@@
G.3 Impact of readout layer depth
@@
G.4 Performance across different graph topologies
In this section, we analyze the models’ performance on different graph topologies in ECHO-Synth. Figures 10, 11 and 12 show respectively diam, ecc and sssp. The results show that although absolute performance varies slightly with the underlying graph topology (e.g., GPS performs better on lobster graphs than on tree-like topologies in SSSP), the relative ranking of models remains consistent, reinforcing the robustness of our findings. Interestingly, we observe that the line topology is consistently the most challenging one. This is expected: in a line graph, every message must pass through a sequence of intermediate nodes, making each node a critical bottleneck for information propagation and amplifying the need for long-range communication.
@@ G.4 Performance across different graph topologies
Appendix H Extended Results
In Table 18, we report additional results on ECHO-Synth benchmark. In particular, we report MAE, MSE and loss (defined as $\log_{10}(\mathrm{MSE})$) obtained on the test set. Similarly, we report the same metrics for ECHO-Charge and ECHO-Energy in Table 19 and Table 20, respectively.
@@ Appendix I Runtimes
Metric Model ECHO-Synth ECHO-Chem
diam sssp ecc ECHO-Charge ECHO-Energy
@@
Training (s) GRIT $2.136_{\pm 0.098}$ $2.898_{\pm 0.663}$ $2.150_{\pm 0.533}$ $55.964_{\pm 0.962}$ $35.246_{\pm 0.539}$
Inference (s) $0.579_{\pm 0.064}$ $0.703_{\pm 0.175}$ $0.632_{\pm 0.132}$ $1.474_{\pm 0.958}$ $1.940_{\pm 0.435}$
MAE $\mathbf{1.014_{\pm 0.046}}$ $\mathbf{0.121_{\pm 0.013}}$ $5.091_{\pm 0.158}$ $7.134_{\pm 6.090}$ $25.508_{\pm 2.507}$
Training (s) PH-DGN $2.840_{\pm 0.060}$ $4.480_{\pm 0.060}$ $3.010_{\pm 0.060}$ $39.405_{\pm 1.753}$ $43.653_{\pm 1.199}$
@@
Appendix J Visualization of GPS Attention Patterns